OpenStudy (anonymous):

Simplify this expression: (4x^5y^3)(3x^8y^7)/6x^4y^10

OpenStudy (ranga):

Group the constant terms, the x terms and the y terms together and then multiply: (4 * 3 / 6) * (x^5 * x^8 / x^4) ( (y^3 * y^7 / y^10) = ?

OpenStudy (ranga):

= (12/6) * x^(5 + 8 - 4) * y^(3 + 7 - 10) = 2 * x^9 * y^0 = 2x^9